OpenStudy (maria1991):

Recall that profit equals revenue minus cost. For the revenue and cost functions​ shown, answer parts ​(a) through​ (e) below. R(x)=440x-2x^2 and c=200x+6120 a) what is break even point? I am stuck in figuring out how the solution is -60 6√15 When I do the quadratic function?

OpenStudy (maria1991):

i know that 120/ 2= 60 but the 6 √15 is where i am confused at.

OpenStudy (maria1991):

the whole quadratic function is 120 √120^2-4(1)(3060) divide by 2

OpenStudy (maria1991):

please help

OpenStudy (maria1991):

do u need more info??

OpenStudy (zpupster):

r-c gives -2x^2 +440x - 200x -6120 or factoring out -2 -2(x^2 - 120x + 3060) = 0 i think this is what you got??

OpenStudy (zpupster):

dividing both sides by -2 gives x^2 - 120x + 3060 = 0 are we ok so far??

OpenStudy (maria1991):

yeah

OpenStudy (maria1991):

when we do the quadratic function, im stuck at figuring out why the answer is 60 6√15

OpenStudy (zpupster):

we can put in a calculator http://www.math.com/students/calculators/source/quadratic.htm or use the quadratic formula i used complete the square x^2 - 120x + 3060 =0 subtracting 3060 to both sides gives x^2 - 120x =-3060 and then add 120/2 and square it gives x^2 - 120x + 3600 =-3060 + 3600 (x-60)^2 = 540

OpenStudy (zpupster):

sqrt bth sides x-60 = + or - sqrt 540

OpenStudy (zpupster):

540 = 36*15

OpenStudy (zpupster):

doyou see it now maria??

OpenStudy (maria1991):

im looking at the way its done over here, i will attach it

OpenStudy (zpupster):

sqrt540 = sqrt6*6*15 so sqrt= 6sqrt 15

OpenStudy (maria1991):

im trying to do an online assignment and it showing me to do it this way

OpenStudy (maria1991):

im just trying to figure out the 6/15 part

OpenStudy (zpupster):

sqrt 2160 =sqrt(144*15) = 12sqrt15 divided by the 2 gives 6sqrt15

OpenStudy (zpupster):

same as what i got completing the square

OpenStudy (zpupster):

can you see it now??

OpenStudy (maria1991):

how u get 144? you jst have to figure out what multiplies to 2160?

OpenStudy (zpupster):

yes

OpenStudy (maria1991):

cause i was doing 2160square root equals 46.48

OpenStudy (maria1991):

I get it now just how you know which factor to use??

OpenStudy (zpupster):

just try squares 4 9 16 ... so on

OpenStudy (maria1991):

ohh sorry it been a while since I learned this stuff

OpenStudy (maria1991):

thanks for your help

OpenStudy (zpupster):

no problem.